OpenStudy (anonymous):

Can someone please help me evaluate the limit below. I am very confused on what to do.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ 1-2\cos x }{ 2x }\]

OpenStudy (anonymous):

@uri could you possibly help?

OpenStudy (freckles):

Well what happens if you plug in 0?

OpenStudy (anonymous):

it becomes 1/0

OpenStudy (anonymous):

therefore we need to go furter to evaluate but im not very good with trig identities so i dont know the next step aka which trig identity i can use

OpenStudy (anonymous):

hello?

OpenStudy (freckles):

nope you don't have to anything else if you had 0/0 we would need go further

OpenStudy (freckles):

and actually you get -1/0 but still this means we are done and the limit is ?????

OpenStudy (anonymous):

undefined?

OpenStudy (freckles):

Yeah Or does not exist or (dne people say for short)

OpenStudy (anonymous):

but that isn't how we learned in class. my teacher told us that we needed to use trig identities when we worked with practice problems.

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x}\] if we had this one we have to go further because we do have 0/0

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x} \cdot \frac{1+\cos(x)}{1+\cos(x)} \\ \lim_{x \rightarrow 0}\frac{1-\cos^2(x)}{x(1+\cos(x)} \\ \lim_{x \rightarrow 0}\frac{\sin^2(x)}{x(1+\cos(x)} \\ \lim_{x \rightarrow 0}\frac{\sin(x)}{x} \frac{\sin(x)}{1} \frac{1}{1+\cos(x)} \\ \lim_{x \rightarrow 0} \frac{\sin(x)}{x} \cdot \lim_{x \rightarrow 0}\sin(x) \cdot \lim_{x \rightarrow 0} \frac{1}{1+\cos(x)} \\ (1)(0)(\frac{1}{1+1}) \\ (1)(0)(\frac{1}{2}) \\ 0 \] But yeah we didn't have that function

OpenStudy (anonymous):

ohhhh, okay. Thank you very much. I'm going to have to keep practicing these to get better. Thank you

OpenStudy (freckles):

if you had 0/0 you would definitely need those trig identities and recall some trig limits

OpenStudy (freckles):

Like the one I mentioned above...\[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1 \text{ and } \lim_{x \rightarrow 0}\frac{1-\cos(x)}{x}\]

OpenStudy (freckles):

So do you think you would have to work further on this one... \[\lim_{x \rightarrow 4}\frac{x+2}{x^2-16}\]?

OpenStudy (freckles):

like if we plug in 4 we would get 6/0 so we know the limit doesn't exist we wouldn't have to bring out more algebra...

OpenStudy (anonymous):

okay thank you so much!